# Emulation Capacity between Idempotent Channels

### v1

@@ 1 Introduction
In this article, we study the interconversion problem for more general channels. We start by considering the zero-error regime. We say that a channel $\mathcal{G}$ can emulate another channel $\mathcal{F}$ with rate $r$, if there exist an encoding and a decoding channel $\mathcal{E}$ and $\mathcal{D}$ (respectively) as well as a pair of integers $(k,n)$ such that
@@
Let $\mathcal{F}=\mathcal{F}\circ\mathcal{F}$ and $\mathcal{G}=\mathcal{G}\circ\mathcal{G}$ be two idempotent channels with $\lambda(\mathcal{F})\neq(1)$ [^1] If $\lambda(\mathcal{F})=(1)$, then $\mathcal{F}$ is a replacer channel (see Proposition 1) and it is simple to see that we can emulate it with any other idempotent channel $\mathcal{G}$ with an arbitrarily large rate., then
@@
Next, we consider the setting beyond the zero-error regime, where the emulation does not need to be exact: the rate $r$ is achievable with error $\delta\in[0,1]$ if there exists channel $\mathcal{E},\mathcal{D}$ and $k,n$ with $r\leq k/n$ such that
@@ 2 Preliminaries and notations
A *-algebra $\mathcal{A}\subseteq\mathcal{L}(\mathcal{H})$ is a $\mathbb{C}$-linear subspace of operators which is stable under the multiplication and and the * operation of $\mathcal{L}(\mathcal{H})$, i.e. taking the transpose conjugate. That is, for all $x,y\in\mathcal{A}$, $xy\in\mathcal{A}$ and for all $x\in\mathcal{A}$, $x^{*}\in\mathcal{A}$. Furthermore, all *-algebras in this work are assumed to be unital, that is there is an identity $\mathds{1}\in\mathcal{A}$ satisfying $\mathds{1}^{*}=\mathds{1}$ and, for all $x\in\mathcal{A}$, $\mathds{1}x=x\mathds{1}=x$. Such an identity operator is necessarily unique in each *-algebra.
@@
$\mathcal{A},\mathcal{B}$ *-algebras
@@
$x^{*}$ Hermitian conjugate of $x$
$\|\cdot\|_{p}$ $\ell_{p}$-norm Ref. eq. (5)
@@
$\operatorname{Rg}(\mathcal{F})$ Range of $\mathcal{F}$ Ref. eq. (6)
$\lambda(\mathcal{F})$ Shape vector of $\mathcal{F}$ Ref. Prop. 1
@@ 3 Decomposition properties of idempotent channels
In this section, we discuss the decomposition properties of the range of idempotent channels, which are summarized in the following propositions. The range of a map $f:\Omega\rightarrow\Omega^{\prime}$, denoted $\operatorname{Rg}(f)$, is the set of all possible images over the domain $\Omega$:
$\operatorname{Rg}(f)=\{f(x)~|~x\in\Omega\}.$ (6)
Proposition 1 (Structure of the range of idempotent channels).
@@
$\mathcal{H}=\bigoplus_{k=1}^{K}\mathcal{H}_{k,1}\otimes\mathcal{H}_{k,2},$ (7)
and a set of density operators $\{\rho_{k}\in\mathcal{D}(\mathcal{H}_{k,2})~|~k\in[K]\}$, such that for all $x\in\mathcal{L}(\mathcal{H})$:
$\mathcal{F}^{*}(x)=\sum_{k=1}^{K}\operatorname{\operatorname{tr}}_{k,2}(P_{k}xP_{k}(\mathds{1}_{d_{k}}\otimes\rho_{k}))\otimes\mathds{1}_{m_{k}},$ (8)
@@
$\mathcal{F}(x)=\sum_{k=1}^{K}\operatorname{\operatorname{tr}}_{k,2}(P_{k}xP_{k})\otimes\rho_{k},$ (9)
where, for all $k\in[K]$, $P_{k}$ is the orthogonal projector onto $\mathcal{H}_{k,1}\otimes\mathcal{H}_{k,2}$, $d_{k}=\dim(\mathcal{H}_{k,1})\geq 1$, $m_{k}=\dim(\mathcal{H}_{k,2})$ and $\operatorname{\operatorname{tr}}_{k,2}:\mathcal{L}(\mathcal{H}_{k,1}\otimes\mathcal{H}_{k,2})\rightarrow\mathcal{L}(\mathcal{H}_{k,1})$ is the partial trace over $\mathcal{H}_{k,2}$, where the argument of this partial trace is seen as an operator in $\mathcal{L}(\mathcal{H}_{k,1}\otimes\mathcal{H}_{k,2})$ by restriction of its input and output to $\mathcal{H}_{k,1}\otimes\mathcal{H}_{k,2}$, which is a subspace of $\mathcal{H}$ for every $k\in[K]$.
Furthermore, the range of $\mathcal{F}^{*}$ is a *-algebra and the following decompositions hold:
$\operatorname{Rg}(\mathcal{F}^{*})=\bigoplus_{k=1}^{K}\mathcal{L}(\mathcal{H}_{k,1})\otimes\mathds{1}_{m_{k}},$ (10)
$\operatorname{Rg}(\mathcal{F})=\bigoplus_{k=1}^{K}\mathcal{L}(\mathcal{H}_{k,1})\otimes\rho_{k}.$ (11)
Note that $\mathcal{F}^{*}$ may not be faithful, indeed the states $\{\rho_{k}\in\mathcal{D}(\mathcal{H}_{k,2})~|~k\in[K]\}$ are not assumed to be full-rank.
@@
where the matrix expression is written for the canonical base of ${\mathbb{C}}^{2}$. Then, in the notations of eq. (7), we have $\mathcal{H}={\mathbb{C}}^{2}$, $K=1$, $\mathcal{H}_{k,1}={\mathbb{C}}$ and $\mathcal{H}_{k,2}={\mathbb{C}}^{2}$ so that $\lambda(\mathcal{F})=(1)$. Note also that, in this case, $\mathcal{F}$ is not faithful as $|0\rangle\!\langle 0|$ is not full-rank in $\mathcal{L}({\mathbb{C}}^{2})$.
@@
We remark that we have $\lambda(\mathcal{F})=(1)$ if and only if $\mathcal{F}$ is a replacer channel, i.e. $\mathcal{F}:x\mapsto\operatorname{\operatorname{tr}}(x)\rho$ with $\rho$ a fixed state. This fact follows easily from eq. (9) taking $K=1$.
@@
[…]
@@
[…]
$\mathcal{F}=\mathcal{D}\circ\widetilde{\mathcal{F}}\circ\mathcal{E}$. (12)

### v2

@@ 1 Introduction
In this article, we study the interconversion problem for more general channels. We start by considering the zero-error regime. We say that a channel $\mathcal{G}$ can emulate another channel $\mathcal{F}$ with rate $r$, if there exists an encoding and a decoding channel $\mathcal{E}$ and $\mathcal{D}$ (respectively) as well as a pair of integers $(k,n)$ such that
@@
Let $\mathcal{F}=\mathcal{F}\circ\mathcal{F}$ and $\mathcal{G}=\mathcal{G}\circ\mathcal{G}$ be two idempotent channels with $\lambda(\mathcal{F})\neq(1)$ [^1] If $\lambda(\mathcal{F})=(1)$, then $\mathcal{F}$ is basically a replacer channel (see Proposition 1) and it is simple to see that we can emulate it with any other idempotent channel $\mathcal{G}$ with an arbitrarily large rate., then
@@
Next, we consider the setting beyond the zero-error regime, where the emulation does not need to be exact: the rate $r$ is achievable with error $\delta\in[0,1]$ if there exists channels $\mathcal{E},\mathcal{D}$ and $k,n$ with $r\leq k/n$ such that
@@ 2 Preliminaries and notations
In this article, we actually consider only *-subalgebras of $\mathcal{L}(\mathcal{H})$. Such a *-subalgebra $\mathcal{A}\subseteq\mathcal{L}(\mathcal{H})$ is a linear subspace of $\mathcal{L}(\mathcal{H})$ which is stable under the multiplication and the * operation of $\mathcal{L}(\mathcal{H})$, i.e. taking the transpose conjugate. That is, for all $x,y\in\mathcal{A}$, $xy\in\mathcal{A}$ and for all $x\in\mathcal{A}$, $x^{*}\in\mathcal{A}$. Furthermore, all *-algebras in this work are assumed to be unital, that is there is an identity $\mathds{1}\in\mathcal{A}$ satisfying $\mathds{1}^{*}=\mathds{1}$ and, for all $x\in\mathcal{A}$, $\mathds{1}x=x\mathds{1}=x$. Such an identity operator is necessarily unique in each *-algebra.
@@
$\mathcal{A},\mathcal{B}$ *-subalgebras of $\mathcal{L}(\mathcal{H})$
@@
$x^{*}$ Hermitian adjoint of $x$
$\|\cdot\|_{p}$ $\ell_{p}$-norm Eq. (5)
@@
$\widehat{\mathcal{F}}$ Reduced channel associated to $\mathcal{F}$ Prop. 1, eq. (9)
$\operatorname{Rg}(\mathcal{F})$ Range of $\mathcal{F}$
$\lambda(\mathcal{F})$ Shape vector of $\mathcal{F}$ Prop. 1
@@ 3 Decomposition properties of idempotent channels
In this section, we discuss the decomposition properties of the range of idempotent channels, which are summarized in the following propositions. The range of a map $f:\Omega\rightarrow\Omega^{\prime}$, denoted $\operatorname{Rg}(f)$, is the set of all possible images over its domain $\Omega$, i.e. $\operatorname{Rg}(f)=\{f(x)~|~x\in\Omega\}$. The range of an idempotent channel coincides with its set of fixed points and the structure of this space is well understood, see for example the lecture notes [19]. The following proposition summarizes the properties that we use.
Proposition 1 (Structure of the range of idempotent channels, see [19]).
@@
$\mathcal{H}=\mathcal{H}_{0}\oplus\mathcal{H}_{1},$ (6)
with $\mathcal{H}_{0}=\operatorname{\operatorname{supp}}(\mathcal{F}(\mathds{1}))$ and $\mathcal{H}_{1}=\ker(\mathcal{F}(\mathds{1}))$, and $\mathcal{H}_{0}$ can further be decomposed as the orthogonal direct sum
$\mathcal{H}_{0}=\bigoplus_{k=1}^{K}\mathcal{H}_{k,1}\otimes\mathcal{H}_{k,2},$ (7)
so that there is a set of density operators $\{\rho_{k}\in\mathcal{D}(\mathcal{H}_{k,2})~|~k\in[K]\}$, such that
$\operatorname{Rg}(\mathcal{F})=\Bigl(\bigoplus_{k=1}^{K}\mathcal{L}(\mathcal{H}_{k,1})\otimes\rho_{k}\Bigr)\oplus 0.$ (8)
Let $e=\mathcal{F}(\mathds{1})^{0}\in\mathcal{L}(\mathcal{H})$ be the orthogonal projector on $\mathcal{H}_{0}$, and $V:\mathcal{H}_{0}\rightarrow\mathcal{H}$ be the isometry embedding $\mathcal{H}_{0}$ into $\mathcal{H}$ such that $VV^{*}=e$ and $V^{*}V=\mathds{1}_{0}$ the identity operator in $\mathcal{L}(\mathcal{H}_{0})$. The reduced quantum channel associated to $\mathcal{F}$, denoted $\widehat{\mathcal{F}}:\mathcal{L}(\mathcal{H}_{0})\rightarrow\mathcal{L}(\mathcal{H}_{0})$ is then defined as
$\widehat{\mathcal{F}}:x\mapsto V^{*}\mathcal{F}(VxV^{*})V,$ (9)
which is idempotent and whose range can be written as
$\operatorname{Rg}(\widehat{\mathcal{F}})=\bigoplus_{k=1}^{K}\mathcal{L}(\mathcal{H}_{k,1})\otimes\rho_{k}.$ (10)
Furthermore, we can express $\widehat{\mathcal{F}}$ and its adjoint $\widehat{\mathcal{F}}^{*}$ in the following way:
$\widehat{\mathcal{F}}(x)=\sum_{k=1}^{K}\operatorname{\operatorname{tr}}_{k,2}(P_{k}xP_{k})\otimes\rho_{k},$ (11)
@@
$\widehat{\mathcal{F}}^{*}(x)=\sum_{k=1}^{K}\operatorname{\operatorname{tr}}_{k,2}(P_{k}xP_{k}(\mathds{1}_{d_{k}}\otimes\rho_{k}))\otimes\mathds{1}_{m_{k}},$ (12)
where, for all $k\in[K]$, $P_{k}$ is the orthogonal projector onto $\mathcal{H}_{k,1}\otimes\mathcal{H}_{k,2}$, $d_{k}=\dim(\mathcal{H}_{k,1})\geq 1$, $m_{k}=\dim(\mathcal{H}_{k,2})$ and $\operatorname{\operatorname{tr}}_{k,2}:\mathcal{L}(\mathcal{H}_{k,1}\otimes\mathcal{H}_{k,2})\rightarrow\mathcal{L}(\mathcal{H}_{k,1})$ is the partial trace over $\mathcal{H}_{k,2}$.
Therefore, the range of $\widehat{\mathcal{F}}^{*}$ is a unital *-subalgebra of $\mathcal{L}(\mathcal{H}_{0})$ and the following decomposition holds:
$\operatorname{Rg}(\widehat{\mathcal{F}}^{*})=\bigoplus_{k=1}^{K}\mathcal{L}(\mathcal{H}_{k,1})\otimes\mathds{1}_{m_{k}}.$ (13)
@@
[…]
$\mathcal{F}(\mathds{1})=\begin{pmatrix}2&0\\ 0&0\end{pmatrix}$.
[…]
$VV^{*}=\begin{pmatrix}1&0\\ 0&0\end{pmatrix}=\mathcal{F}(\mathds{1})^{0}.$
[…]
$\widehat{\mathcal{F}}(x)=V^{*}\mathcal{F}(VxV^{*})V=x,$
so that $\widehat{\mathcal{F}}$ is the identity on $\mathcal{L}({\mathbb{C}})$.
@@
Furthermore, we remark that we have $\lambda(\mathcal{F})=(1)$ if and only if its reduced channel is a replacer channel, i.e. $\widehat{\mathcal{F}}:x\mapsto\operatorname{\operatorname{tr}}(x)\rho$ with $\rho$ a fixed state. This fact follows easily from eq. (11) taking $K=1$.
Before we present the proof of Proposition 1, we need to state the two following lemmas which will be used in the proof.
Lemma 3.
Let $\Phi:\mathcal{L}(\mathcal{H})\rightarrow\mathcal{L}(\mathcal{H})$ be a positive linear map [^2] Here ’positive’ means that $x\geq 0$ implies $\Phi(x)\geq 0$., let $\mathcal{H}_{0}=\operatorname{\operatorname{supp}}(\Phi(\mathds{1}))$ and let $V:\mathcal{H}_{0}\rightarrow\mathcal{H}$ be an isometry embedding $\mathcal{H}_{0}$ into $\mathcal{H}$ so that $VV^{*}\in\mathcal{L}(\mathcal{H})$ is the orthogonal projection on $\mathcal{H}_{0}$. Then for all $x\in\mathcal{L}(\mathcal{H})$,
$VV^{*}\Phi(x)VV^{*}=\Phi(x).$
Proof.
Let $x\in\mathcal{L}(\mathcal{H})$ be positive semidefinite, then
$0\leq x\leq\|x\|\mathds{1},$
so that by positivity of $\Phi$, we have
$0\leq\Phi(x)\leq\|x\|\Phi(\mathds{1}).$
Hence, $\Phi(x)$ is supported on the support of $\Phi(\mathds{1})$ which is, by definition, $\mathcal{H}_{0}$. Furthermore, as $x$ and $\mathds{1}$ are positive semidefinite, $\Phi(x)$ and $\Phi(\mathds{1})$ are positive semidefinite as well so that their respective range (as linear maps on $\mathcal{H}$) coincides with their support. Hence, we have
$VV^{*}\Phi(x)VV^{*}=\Phi(x).$
Now, if $x\in\mathcal{L}(\mathcal{H})$ is arbitrary, we can use the polarisation identity (see eq. (1.45) in [19]) to write
$x=\frac{1}{4}\sum_{k=0}^{3}i^{k}(x+i^{k}\mathds{1})^{*}(x+i^{k}\mathds{1}),$
so that $x$ is a linear combination of four positive semidefinite elements of $\mathcal{L}(\mathcal{H})$. Now, by linearity, we get
$VV^{*}\Phi(x)VV^{*}=\Phi(x).$
Furthermore, we remark that the set of fixed points and the range of an idempotent function coincide.
Lemma 4.
$\operatorname{Rg}(f)=\{x\in\Omega~|~f(x)=x\}.$ (14)
Proof.
If $x$ is a fixed point, $x=f(x)\in\operatorname{Rg}(f)$. For the other inclusion, if $y=f(x)$, we also have $f(y)=f(f(x))=f(x)=y$, so $y$ is a fixed point. ∎
Proposition 1 then basically follows from statements in Section 6.4 [19] and we detail it here only for the convenience of the reader. Similar derivation can be found in [1] in the Heisenberg picture.
Proof of Proposition 1.
We will first prove that the reduced channel $\widehat{\mathcal{F}}$, defined in eq. (9), is actually a channel. It is clearly completely positive as it is a composition of completely positive maps and it is trace preserving because for all $x\in\mathcal{L}(\mathcal{H}_{0})$,
$\operatorname{\operatorname{tr}}(\widehat{\mathcal{F}}(x))=\operatorname{\operatorname{tr}}(V^{*}\mathcal{F}(VxV^{*})V)=\operatorname{\operatorname{tr}}(VV^{*}\mathcal{F}(VxV^{*})VV^{*})=\operatorname{\operatorname{tr}}(\mathcal{F}(VxV^{*}))=\operatorname{\operatorname{tr}}(VxV^{*})=\operatorname{\operatorname{tr}}(x),$
where the third equality follows from Lemma 3. Furthermore, $\widehat{\mathcal{F}}$ is idempotent as, for $x\in\mathcal{L}(\mathcal{H}_{0})$,
$\widehat{\mathcal{F}}(\widehat{\mathcal{F}}(x))$ $=V^{*}\mathcal{F}(VV^{*}\mathcal{F}(VxV^{*})VV^{*})V$
$=V^{*}\mathcal{F}(\mathcal{F}(VxV^{*}))V$
$=V^{*}\mathcal{F}(VxV^{*})V$
$=\widehat{\mathcal{F}}(x).$
The second equation follows from Lemma 3, and the third from the idempotence of $\mathcal{F}$.
Furthermore, $\widehat{\mathcal{F}}$ admits a full-rank fixed point. Indeed, $\widehat{\mathcal{F}}(V^{*}\mathcal{F}(\mathds{1})V)$ is a fixed point of $\widehat{\mathcal{F}}$, since $\widehat{\mathcal{F}}$ is idempotent, and we have
$\widehat{\mathcal{F}}(V^{*}\mathcal{F}(\mathds{1})V)=V^{*}\mathcal{F}(VV^{*}\mathcal{F}(\mathds{1})VV^{*})V=V^{*}\mathcal{F}\mathcal{F}(\mathds{1})V=V^{*}\mathcal{F}(\mathds{1})V,$
which is full-rank in $\mathcal{H}_{0}=\operatorname{\operatorname{supp}}(\mathcal{F}(\mathds{1}))$, again using Lemma 3 for the second equality.
Therefore, we can apply Theorem 6.12 in [19] to the adjoint $\widehat{\mathcal{F}}^{*}$ of $\widehat{\mathcal{F}}$. Hence, the set of fixed points of $\widehat{\mathcal{F}}^{*}$ is a *-subalgebra of $\mathcal{L}(\mathcal{H}_{0})$. By Lemma 4, since $\widehat{\mathcal{F}}^{*}$ is idempotent, its set of fixed points coincides with its range, so that its range is a *-subalgebra of $\mathcal{L}(\mathcal{H}_{0})$. It is well known (see e.g. [17]), that if $\mathcal{A}\subseteq\mathcal{L}(\mathcal{H}_{0})$ is a *-subalgebra of $\mathcal{L}(\mathcal{H}_{0})$, we can find a basis of $\mathcal{H}_{0}$ in which it can be expressed as an orthogonal direct sum as in eq. (13), which proves this equation. Furthermore, we get by the Proposition 1.5 of [19] that $\widehat{\mathcal{F}}^{*}$ takes the form of eq. (12). Then, taking the adjoint of this equation, we obtain that $\widehat{\mathcal{F}}$$\widehat{\mathcal{F}}^{**}$ can be written as in eq. (11). Then, for $x\in\mathcal{L}(\mathcal{H}_{0})$, we have that
$\widehat{\mathcal{F}}(x)=\sum_{k=1}^{K}\operatorname{\operatorname{tr}}_{k,2}(P_{k}xP_{k})\otimes\rho_{k}\in\bigoplus_{k=1}^{K}\mathcal{L}(\mathcal{H}_{k,1})\otimes\rho_{k}.$
Conversely, if
$x=\sum_{k=1}^{K}x_{k}\otimes\rho_{k}\in\bigoplus_{k=1}^{K}\mathcal{L}(\mathcal{H}_{k,1})\otimes\rho_{k},$
[…]
Finally, to prove eq. (8), let $x\in\operatorname{Rg}(\mathcal{F})$. We then have
$x=\mathcal{F}(x)=VV^{*}\mathcal{F}(x)VV^{*}=VV^{*}xVV^{*},$ (15)
so that $x\in\mathcal{L}(\mathcal{H}_{0})\oplus 0$ where the 0 operator acts on $\mathcal{H}_{1}$. Then, $V^{*}xV\in\mathcal{L}(\mathcal{H}_{0})$ and
$\widehat{\mathcal{F}}(V^{*}xV)=V^{*}\mathcal{F}(VV^{*}xVV^{*})V=V^{*}\mathcal{F}(x)V=V^{*}xV,$
by Lemma 3 and by eq. (15). Hence,
$V^{*}xV\in\bigoplus_{k=1}^{K}\mathcal{L}(\mathcal{H}_{k,1})\otimes\rho_{k},$
by eq. (10) and
$x=VV^{*}xVV^{*}\in\bigoplus_{k=1}^{K}\mathcal{L}(\mathcal{H}_{k,1})\otimes\rho_{k}\oplus 0.$
Conversely, if
$x=\begin{pmatrix}x_{0}&0\\ 0&0\end{pmatrix}$,
with
$x_{0}\in\bigoplus_{k=1}^{K}\mathcal{L}(\mathcal{H}_{k,1})\otimes\rho_{k},$
then
$x=Vx_{0}V^{*}=V\widehat{\mathcal{F}}(x_{0})V^{*}=VV^{*}\mathcal{F}(Vx_{0}V^{*})VV^{*}=\mathcal{F}(Vx_{0}V^{*})=\mathcal{F}(x),$
so that $x\in\operatorname{Rg}(\mathcal{F})$. ∎
We will now show that if $\mathcal{F}$ and $\mathcal{G}$ are two idempotent channels, it is equivalent to study the emulation of $\mathcal{F}$ by $\mathcal{G}$ and the emulation of their reduced version, i.e. $\widehat{\mathcal{F}}$ by $\widehat{\mathcal{G}}$. First, we can explicitly find an encoder and a decoder channel to convert any idempotent channel $\mathcal{F}$ into its reduced version $\widehat{\mathcal{F}}$ and vice versa as shown in the following proposition.
Proposition 2.
[…]
Proof.
Let $V:\mathcal{H}_{0}\rightarrow\mathcal{H}$ be the isometry embedding $\mathcal{H}_{0}$ into $\mathcal{H}$ which defines $\widehat{\mathcal{F}}$ as in Proposition 1. For the first equality, let $\widehat{\mathcal{E}}:\mathcal{L}(\mathcal{H})\rightarrow\mathcal{L}(\mathcal{H}_{0})$ be defined by
$\widehat{\mathcal{E}}:x\mapsto V^{*}\mathcal{F}(x)V,$
and $\widehat{\mathcal{D}}:\mathcal{L}(\mathcal{H}_{0})\rightarrow\mathcal{L}(\mathcal{H})$ be defined by
$\widehat{\mathcal{D}}:x\mapsto VxV^{*}.$
Both $\widehat{\mathcal{E}}$ and $\widehat{\mathcal{D}}$ are completely positive trace preserving map. The complete positivity is easy. $\widehat{\mathcal{D}}$ is trace preserving because $\operatorname{\operatorname{tr}}(VxV^{*})=\operatorname{\operatorname{tr}}(V^{*}Vx)=\operatorname{\operatorname{tr}}(x)$ and $\widehat{\mathcal{E}}$ is trace preserving as well because for all $x\in\mathcal{L}(\mathcal{H})$,
$\operatorname{\operatorname{tr}}(V^{*}\mathcal{F}(x)V)=\operatorname{\operatorname{tr}}(VV^{*}\mathcal{F}(x)VV^{*})=\operatorname{\operatorname{tr}}(\mathcal{F}(x))=\operatorname{\operatorname{tr}}(x),$
where the second equality follows from Lemma 3. Then we have the factorisation
$\mathcal{F}=\widehat{\mathcal{D}}\widehat{\mathcal{F}}\widehat{\mathcal{E}}$.
Indeed, let $x\in\mathcal{L}(\mathcal{H})$, we have
$\widehat{\mathcal{D}}\widehat{\mathcal{F}}\widehat{\mathcal{E}}(x)$ $=VV^{*}\mathcal{F}(VV^{*}\mathcal{F}(x)VV^{*})VV^{*}$
$=\mathcal{F}(\mathcal{F}(x))$
$=\mathcal{F}(x),$
where the second equality follows from Lemma 3. Now, for the factorisation of $\widehat{\mathcal{F}}$, let $\mathcal{E}=\widehat{\mathcal{D}}$ and $\mathcal{D}=\widehat{\mathcal{E}}$, for $x\in\mathcal{L}(\mathcal{H}_{0})$, we have:
$\mathcal{D}\mathcal{F}\mathcal{E}(x)$ $=\widehat{\mathcal{E}}\mathcal{F}\widehat{\mathcal{D}}(x)$
$=V^{*}\mathcal{F}(\mathcal{F}(VxV^{*}))V$
$=V^{*}\mathcal{F}(VxV^{*})V$
$=\widehat{\mathcal{F}}(x).$
As a consequence of this proposition, we get that studying the interconversion of two idempotent channels $\mathcal{F}$ and $\mathcal{G}$, is equivalent to study the interconversion of their reduced versions $\widehat{\mathcal{F}}$ and $\widehat{\mathcal{G}}$.
Proposition 3.
[…]
$\inf_{\mathcal{D}\!,\mathcal{E}}\|\mathcal{F}-\mathcal{D}\mathcal{G}\mathcal{E}\|_{\diamond}=\inf_{\mathcal{D}\!,\mathcal{E}}\|\widehat{\mathcal{F}}-\mathcal{D}\widehat{\mathcal{G}}\mathcal{E}\|_{\diamond}$, (17)
where the infimum are taken over encoders and decoders channels $\mathcal{E}$ and $\mathcal{D}$ between appropriate spaces.
Proof.
Let $\mathcal{E}:\mathcal{L}(\mathcal{H}_{\mathcal{F}})\rightarrow\mathcal{L}(\mathcal{H}_{\mathcal{G}})$ and $\mathcal{D}:\mathcal{L}(\mathcal{H}_{\mathcal{G}})\rightarrow\mathcal{L}(\mathcal{H}_{\mathcal{F}})$ be a pair of channels. Let $\mathcal{E}_{1}$, $\mathcal{D}_{1}$, $\mathcal{E}_{2}$, $\mathcal{D}_{2}$ be channels such that
[…]
which exist by Proposition 2. Then,
$\|\mathcal{F}-\mathcal{D}\mathcal{G}\mathcal{E}\|_{\diamond}$ $=\|\mathcal{F}-\mathcal{D}\mathcal{D}_{2}\widehat{\mathcal{G}}\mathcal{E}_{2}\mathcal{E}\|_{\diamond}$
$=\|\mathcal{D}_{1}\|_{\diamond}\|\mathcal{F}-\mathcal{D}\mathcal{D}_{2}\widehat{\mathcal{G}}\mathcal{E}_{2}\mathcal{E}\|_{\diamond}\|\mathcal{E}_{1}\|_{\diamond}$
$\geq\|\mathcal{D}_{1}\mathcal{F}\mathcal{E}_{1}-\mathcal{D}_{1}\mathcal{D}\mathcal{D}_{2}\widehat{\mathcal{G}}\mathcal{E}_{2}\mathcal{E}\mathcal{E}_{1}\|_{\diamond}$
$=\|\widehat{\mathcal{F}}-\mathcal{D}_{1}\mathcal{D}\mathcal{D}_{2}\widehat{\mathcal{G}}\mathcal{E}_{2}\mathcal{E}\mathcal{E}_{1}\|_{\diamond},$
where we used in the second equality that as $\mathcal{E}_{1}$, $\mathcal{D}_{1}$ are channels, $\|\mathcal{E}_{1}\|_{\diamond}=\|\mathcal{D}_{1}\|_{\diamond}=1$.
Therefore,
$\inf_{\mathcal{D}\!,\mathcal{E}}\|\widehat{\mathcal{F}}-\mathcal{D}\widehat{\mathcal{G}}\mathcal{E}\|_{\diamond}\leq\inf_{\mathcal{D}\!,\mathcal{E}}\|\mathcal{F}-\mathcal{D}\mathcal{G}\mathcal{E}\|_{\diamond}.$
We can apply the same line of argument but converting $\widehat{\mathcal{F}}$ to $\mathcal{F}$ and $\mathcal{G}$ to $\widehat{\mathcal{G}}$ to prove the converse inequality. ∎
@@
We now show that the range and the shape vector of idempotent channels preserve the tensor product, a property that will be used in the proofs of our Theorems 1 and 2. Indeed, it will allow us to reduce the study of the problem of emulating $k$ copies of $\mathcal{F}$ using $n$ copies of $\mathcal{G}$ for general pairs of integers $(k,n)$ to the case where $k=n$.
@@
Let $\mathcal{F}_{1}:\mathcal{L}(\mathcal{H}_{1})\rightarrow\mathcal{L}(\mathcal{H}_{1})$ and $\mathcal{F}_{2}:\mathcal{L}(\mathcal{H}_{2})\rightarrow\mathcal{L}(\mathcal{H}_{2})$ be two idempotent channels, we have
$\operatorname{Rg}(\mathcal{F}_{1}\otimes\mathcal{F}_{2})$ $=\operatorname{Rg}(\mathcal{F}_{1})\otimes\operatorname{Rg}(\mathcal{F}_{2}),$ (19)
$\operatorname{Rg}(\mathcal{F}_{1}^{*}\otimes\mathcal{F}_{2}^{*})$ $=\operatorname{Rg}(\mathcal{F}_{1}^{*})\otimes\operatorname{Rg}(\mathcal{F}_{2}^{*}),$ (20)
$\widehat{\mathcal{F}_{1}\otimes\mathcal{F}_{2}}$ $=\widehat{\mathcal{F}_{1}}\otimes\widehat{\mathcal{F}_{2}}$, (21)
$\lambda(\mathcal{F}_{1}\otimes\mathcal{F}_{2})$ $=\lambda(\mathcal{F}_{1})\otimes\lambda(\mathcal{F}_{2}),$ (22)
where $\widehat{\mathcal{F}_{1}\otimes\mathcal{F}_{2}}$, $\widehat{\mathcal{F}}_{2}$ and $\widehat{\mathcal{F}}_{2}$ are respectively the reduced channels associated to $\mathcal{F}_{1}\otimes\mathcal{F}_{2}$, $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$.